OpenStudy (anonymous):

If you have a triangle with the lengths of 36, 60, and 48 what is the sine x?

OpenStudy (anonymous):

where is x?

Parth (parthkohli):

^^ Exactly, can you draw and mark x?

OpenStudy (anonymous):

I will attach a .jpg

Parth (parthkohli):

If anyone doesn't know, \[\LARGE \sin = {opp \over hyp}\]

OpenStudy (anonymous):

your triangle has sides that are a 3 - 4 - 5 right triangle, only each number has been multiplied by 12 so your answer will be either \(\frac{3}{5}\) or \(\frac{4}{5}\) depending on the length of the opposite side

OpenStudy (anonymous):

You have to put it into simple fraction form.

OpenStudy (anonymous):

what parth said. opposite over hypotenuse

Parth (parthkohli):

\[\LARGE \sin(x) = {48 \over 60} = {4 \over 5}\]

Parth (parthkohli):

If you want it in decimals too, it's 0.8 or in percentage it's 80%

OpenStudy (anonymous):

i have another sine cosine and tangent question. For triangle QPS what is the sine cosine and tangent of this triangle

Parth (parthkohli):

Which angle please?

OpenStudy (anonymous):

again its in fraction form. Angle P

Parth (parthkohli):

\[\LARGE \sin(p) = {5 \over 14} \] \[\LARGE \tan(p) = {opposite \over adjacent} = {5 \over 9}\]

OpenStudy (anonymous):

I have a couple more questions if your willing!

Parth (parthkohli):

I am., continue.

OpenStudy (anonymous):

The length of the hypotenuse of a 30 degree 60 degree 90 degree triangle is 28 m. Find the length of the side opposite the 30 degree angle.

Parth (parthkohli):